# DiVE-k: differential visual reasoning for fine-grained image recognition

### v1

@@ 1 Introduction
We explore the task of zero-shot fine-grained image recognition, as a visual reasoning task building on available Large Vision Language Models (LVLMs). Such capabilities are crucial for the generalization of vision systems. In early zero-shot image recognition works, such as in CLIP (radford2021learning), the visual embedding from an image is matched against the text embedding of class names to determine the most likely label. LVLMs, such as QWEN2-VL (wang2024qwen2), contain a Large Language Model (LLM) in themselves and are able to use their vast language knowledge with unified multimodal pre-training to achieve impressive capabilities in zero-shot recognition. however, the accuracy for fine-grained recognition is limited. We aim to improve accuracy by fine-tuning on a subset of categories of a new dataset (called the “base” set) and test on “novel” categories for which no training examples are seen. This setting is common (zhou2022learning) and relevant for adapting models to new domains with limited training data.
@@
Inspired by the success of Chain of Thought (CoT) reasoning (wei2022chain) using Reinforcement Learning (RL) (guo2025deepseek) for pure language tasks such as mathematics and coding (shao2024deepseekmath; jiang2023mistral7b), ViRFT (liu2025visual) extended this idea to vision tasks such as image classification. The key idea in ViRFT is to construct a verifiable reward (lambert2024tulu) for image classification which allows visual reasoning RL training via GRPO (shao2024deepseekmath). However, their verifiable reward obtained through exact string match between the category name and the model’s final answer is brittle: (i) requiring ad hoc string post-processing and model responses may use scientific or common names that cannot be validated by string matching; (ii) encourages memorization of the training category names (Section 4.2.1); and (iii) fails to incentivize attribute-level, discriminative reasoning see Figure 1b. This leads to a weak base to novel generalization and a tendency to ignore useful text knowledge when visual evidences alone are ambiguous.
@@ 2 Prior Work
Zero-shot fine-grained image classification Vision Language Models (VLMs) (radford2021learning; li2022blip; tschannen2025siglip; yuan2021florence) use the idea of aligning image with text to achieve zero-shot learning (lampert2013attribute; socher2013zero; wang2018zero; zhang2017learning). Although very competent for image-text alignment, these models have limited world knowledge unlike LLMs (radford2019language; brown2020language; touvron2023llama; chowdhery2023palm). An early line of research proposes the idea of prompt learning (zhou2022conditional; khattak2023self) where a prompt vector is learned for text prompt’s context words. zheng2024large uses LLM’s knowledge to learn the prompt vector. Another approach to bridge the knowledge gap is by combining the perceptual strengths of VLMs with the linguistic abilities of LLM (esfandiarpoor2023follow; menon2022visual; pratt2023does; zeng2022socratic; novack2023chils; roth2023waffling). Our work in part is inspired by FuDD (esfandiarpoor2023follow) that uses a multi-stage reasoning by using LLM knowledge to find pairwise discriminative features to compliment VLM. However, FuDD generates a fixed set of text prompts offline, limiting its ability to adapt its reasoning strategy to the specific difficulty of each input. Other related work tries to use LLM as a tool for reasoning through programming and language reasoner (chen2023see; gupta2023visual; zhang2023multimodal; suris2023vipergpt). While these methods improved model performance, the separation between the vision and language modules creates a bottleneck, inhibiting seamless and integrated reasoning across modalities (liu2023visual; liu2024improved).
@@
Recent LVLMs (bai2025qwen2; hurst2024gpt; team2023gemini; team2025gemma) have excellent visual understanding, such as VQA (shao2024visual), combining visual encoding directly into a LLM architecture. This opens up the possibility of joint vision and text reasoning capability (team2025kimi_vl; bai2025qwen2; team2025gemma). Additionally, recent success of RL for reasoning on maths and coding tasks (jaech2024openai; guo2025deepseek; team2025kimi) have transformed the post-training reasoning research and there is a growing interest to extend these ideas to LVLMs.
@@
Enhancing Reasoning with Reinforcement Learning . Building upon the seminal work of In-context learning (brown2020language) and CoT (wei2022chain), the field has moved beyond static prompting by applying RL to fine-tune these reasoning processes (jaech2024openai; guo2025deepseek; team2025kimi). By treating the generation of a reasoning chain as a sequential decision-making problem, RL-based methods can train models to produce more accurate and explainable solutions. Recent breakthrough in DeepSeek-R1 (shao2024deepseekmath) showed the effectiveness of CoT based training further making it more efficient using their GRPO algorithm. Inspired by these success in text-based reasoning, a growing body of work apply RL to enhance the reasoning capabilities of LVLMs for vision-centric tasks, such as image classification (liu2025visual; li2025think), Object detection, Grounding (liu2025visual; shen2025vlm), and Visual Question Answering (cao2025ground; sarch2025grounded; fan2025grit). Within our target domain of fine-grained image recognition, the most pertinent work is ViRFT (liu2025visual), which trains an LVLM using exact string matching reward to foster visual reasoning. However, this reward mechanism proves brittle, failing to generalize in base-to-novel settings. Furthermore, it does not fully leverage the LLM’s inherent knowledge about fine-grained categories to incentivize the differential reasoning necessary for distinguishing among similar options. zhu2025surprising; chen2025pass proposes to enhance LLM performance by improving Pass@k accuracy. In contrast, we use model’s own knowledge of Pass@k to improve their reasoning, specifically for vision-language task.
@@ 4.1 Experimental Setup
Baselines and Datasets. We compare our results against the various baselines: 1) pre-trained QWEN2.5-VL-7B (bai2025qwen2) 2) Supervised Fine Tuning (SFT) (zheng2024llamafactory) using full model and LoRA (hu2022lora) training. 3) ViRFT (liu2025visual) trained under the same base-novel setting. 4) QWEN2.5-VL-7B and ViRFT using our inference pipeline 5) Consistency as accuracy: During the two-step inference process, use the most consistent prediction from $K$ generation as final prediction. We also report performance of the proprietary models, Gemini2.5-flash-light(comanici2025gemini), GPT-5-mini (gpt5systemcard) and Grok4-fast (grok4fast), with closed weights accessible through APIs (openrouter).
We evaluate our method on five standard fine-grained image classification dataset across various domains, including OxfordFlowers-102 (Nilsback08), CUB-200 (wah2011caltech), OxfordPets-37 (parkhi12a), StanfordCars-196 (krause20133d) and FGVC Aircraft-100 (maji2013fine). For the base-novel split, we follow previous work (Khattak_2023_ICCV) to divide the categories into equal halves for base and novel. For instance in CUB-200, the first 100 categories are considered base and the other 100 are considered novel and similarly for other datasets. Note that neither images nor the category names of the novel classes are seen during training.
@@
Evaluation setting and Metric. We evaluate our method under two distinct zero-shot settings. The first is the standard zero-shot base-to-novel generalization, following (Zheng_2024_CVPR), where a separate model is trained on the base classes of each dataset. The second is our proposed mixed-dataset setting, designed to assess cross-domain generalization capabilities. Here, a single model is trained on a unified dataset constructed by combining the base classes from all datasets. Additionally, we also evaluate our method under few-shot classification setting with 4 shots per class. For performance measurement, we report the classification accuracy on base and novel classes, along with their Harmonic Mean (HM).
For evaluation, we use the LLM gemini-2.5-flash-lite (comanici2025gemini), where we provide the ground-truth category name and model predicted category name and ask the LLM if they belong to the same fine-grained category or not. This specifically helps us evaluate better for the answers where model responds a scientific name and the provided ground-truth is common name and vice-versa. We provide more details in Appendix A.1.2
Implementation Details. We use the pre-trained Qwen2.5-VL-7B-Instruct (bai2025qwen2) as our base model and perform the RL training using the proposed method. We use K=20 during offline option generation using $K$ rollouts and other details are provided in Appendix A.1.1. For Zero Shot Base-to-Novel training, we train the model for 400 steps, whereas for mixed data training, we train it for 1 epoch. For few-shot training, we train each model for 200 steps following ViRFT. All models are trained on three A6000 GPUs with 48GB memory with an overall batch size of 6. For GRPO, a total of 4 responses are generated for each input sample. Following ViRFT, we use learning rate of $10^{-6}$, AdamW optimizer, linear scheduler and set $\lambda_{m}=\lambda_{f}=1$. For ViRFT training, we use their official code from github with some modification discussed in Appendix A.1.1.
@@ 4.2.1 Zero-Shot Base-to-Novel Generalization
We also evaluate supervised fine-tuning (SFT) as a baseline. Although SFT yields strong accuracy gains on base categories, its performance deteriorates sharply on novel categories, with an average accuracy drop of 33.5% relative to the base model and a 19.9% reduction in HM. This sharp degradation highlights SFT’s inability to generalize under the base-to-novel transfer setting. To provide a broader perspective, we also include results from proprietary models which are likely much larger. In this context, our method surpasses Grok4-fast and is on par with Gemini2.5-flash-light, though GPT-5-mini leads. We provide more results using Gemma-3 (team2025gemma) base model in Appendix A.3.
@@ 4.3 Ablation Studies
Our ablation into the MCQ option generation reveals that the option construction strategy is critical for model’s performance. As detailed in Table 3, randomly selecting categories proved suboptimal, yielding only marginal gains and failing to instill robust reasoning capabilities. While employing a text embedding model from gemini embeddings (lee2025gemini) to generate semantically similar options offered some improvement, our proposed top-k as options proves significantly more effective. By sampling options directly from the same base model’s top-k generations achieved a substantial classification accuracy gain of 12% on base classes and 6.9% on novel classes compared to base model. This demonstrates that leveraging the base model’s own knowledge distribution to generate options is the optimal strategy for training, leading to superior generalization on both base and novel sets.
@@ A.1.2 Evaluation details
In Figure 8, we provide the prompt used for second step of multiple choice question. Here {options} refers to the options obtained from the first step of our pipeline. In Figure 9, we provide the prompt used during the evaluation. {groundtruth} refers to the groundtruth category name and {prediction} refers to the model’s predicted answer. We use "google/gemini-2.5-flash-lite-preview-06-17" as LLM for evaluation from openrouter (openrouter) API.

### v2

@@ 1 Introduction
We explore the task of zero-shot fine-grained image recognition, as a visual reasoning task building on available Large Vision Language Models (LVLMs). Such capabilities are crucial for the generalization of vision systems. In early zero-shot image recognition works, such as in CLIP (Radford et al., 2021), the visual embedding from an image is matched against the text embedding of class names to determine the most likely label. LVLMs, such as QWEN2-VL (Wang et al., 2024), contain a Large Language Model (LLM) in themselves and are able to use their vast language knowledge with unified multimodal pre-training to achieve impressive capabilities in zero-shot recognition. however, the accuracy for fine-grained recognition is limited. We aim to improve accuracy by fine-tuning on a subset of categories of a new dataset (called the “base” set) and test on “novel” categories for which no training examples are seen. This setting is common (Zhou et al., 2022b) and relevant for adapting models to new domains with limited training data.
@@
Inspired by the success of Chain of Thought (CoT) reasoning (Wei et al., 2022) using Reinforcement Learning (RL) (Guo et al., 2025) for pure language tasks such as mathematics and coding (Shao et al., 2024b; Jiang et al., 2023), ViRFT (Liu et al., 2025) extended this idea to vision tasks such as image classification. The key idea in ViRFT is to construct a verifiable reward (Lambert et al., 2024) for image classification which allows visual reasoning RL training via GRPO (Shao et al., 2024b). However, their verifiable reward obtained through exact string match between the category name and the model’s final answer is brittle: (i) requiring ad hoc string post-processing and model responses may use scientific or common names that cannot be validated by string matching; (ii) encourages memorization of the training category names (Section 4.2.1); and (iii) fails to incentivize attribute-level, discriminative reasoning see Figure 1b. This leads to a weak base to novel generalization and a tendency to ignore useful text knowledge when visual evidences alone are ambiguous.
@@ 2 Prior Work
Zero-shot fine-grained image classification Vision Language Models (VLMs) (Radford et al., 2021; Li et al., 2022; Tschannen et al., 2025; Yuan et al., 2021) use the idea of aligning image with text to achieve zero-shot learning (Lampert et al., 2013; Socher et al., 2013; Wang et al., 2018; Zhang et al., 2017). Although very competent for image-text alignment, these models have limited world knowledge unlike LLMs (Radford et al., 2019; Brown et al., 2020; Touvron et al., 2023; Chowdhery et al., 2023). An early line of research proposes the idea of prompt learning (Zhou et al., 2022a; Khattak et al., 2023b) where a prompt vector is learned for text prompt’s context words. Zheng et al. (2024c) uses LLM’s knowledge to learn the prompt vector. Another approach to bridge the knowledge gap is by combining the perceptual strengths of VLMs with the linguistic abilities of LLM (Esfandiarpoor and Bach, 2023; Menon and Vondrick, 2022; Pratt et al., 2023; Zeng et al., 2022; Novack et al., 2023; Roth et al., 2023). Our work in part is inspired by FuDD (Esfandiarpoor and Bach, 2023) that uses a multi-stage reasoning by using LLM knowledge to find pairwise discriminative features to compliment VLM. However, FuDD generates a fixed set of text prompts offline, limiting its ability to adapt its reasoning strategy to the specific difficulty of each input. Other related work tries to use LLM as a tool for reasoning through programming and language reasoner (Chen et al., 2023; Gupta and Kembhavi, 2023; Zhang et al., 2023; Surís et al., 2023). While these methods improved model performance, the separation between the vision and language modules creates a bottleneck, inhibiting seamless and integrated reasoning across modalities (Liu et al., 2023; 2024).
@@
Recent LVLMs (Bai et al., 2025; Hurst et al., 2024; Team et al., 2023; 2025a) have excellent visual understanding, such as VQA (Shao et al., 2024a), combining visual encoding directly into a LLM architecture. This opens up the possibility of joint vision and text reasoning capability (Team et al., 2025c; Bai et al., 2025; Team et al., 2025a). Additionally, recent success of RL for reasoning on maths and coding tasks (Jaech et al., 2024; Guo et al., 2025; Team et al., 2025b) have transformed the post-training reasoning research and there is a growing interest to extend these ideas to LVLMs.
@@
Enhancing Reasoning with Reinforcement Learning . Building upon the seminal work of In-context learning (Brown et al., 2020) and CoT (Wei et al., 2022), the field has moved beyond static prompting by applying RL to fine-tune these reasoning processes (Jaech et al., 2024; Guo et al., 2025; Team et al., 2025b). By treating the generation of a reasoning chain as a sequential decision-making problem, RL-based methods can train models to produce more accurate and explainable solutions. Recent breakthrough in DeepSeek-R1 (Shao et al., 2024b) showed the effectiveness of CoT based training further making it more efficient using their GRPO algorithm. Inspired by these success in text-based reasoning, a growing body of work apply RL to enhance the reasoning capabilities of LVLMs for vision-centric tasks, such as image classification (Liu et al., 2025; Li et al., 2025), Object detection, Grounding (Liu et al., 2025; Shen et al., 2025), and Visual Question Answering (Cao et al., 2025; Sarch et al., 2025; Fan et al., 2025). Within our target domain of fine-grained image recognition, the most pertinent work is ViRFT (Liu et al., 2025), which trains an LVLM using exact string matching reward to foster visual reasoning. However, this reward mechanism proves brittle, failing to generalize in base-to-novel settings. Furthermore, it does not fully leverage the LLM’s inherent knowledge about fine-grained categories to incentivize the differential reasoning necessary for distinguishing among similar options. Zhu et al. (2025); Chen et al. (2025) proposes to enhance LLM performance by improving Pass@k accuracy. In contrast, we use model’s own knowledge of Pass@k to improve their reasoning, specifically for vision-language task.
@@ 4.1 Experimental Setup
Baselines and Datasets. We compare our results against the various baselines: 1) pre-trained QWEN2.5-VL-7B (Bai et al., 2025) 2) Supervised Fine Tuning (SFT) (Zheng et al., 2024a) using full model and LoRA (Hu et al., 2022) training. 3) ViRFT (Liu et al., 2025) trained under the same base-novel setting. 4) QWEN2.5-VL-7B and ViRFT using our inference pipeline 5) Consistency as accuracy: During the two-step inference process, use the most consistent prediction from $K$ generation as final prediction. We also report performance of the proprietary models, Gemini2.5-flash-light(Comanici et al., 2025), GPT-5-mini (OpenAI, 2025) and Grok4-fast (xAI, 2025), with closed weights accessible through APIs (OpenRouter, 2024).
We evaluate our method on five standard fine-grained image classification dataset across various domains, including OxfordFlowers-102 (Nilsback and Zisserman, 2008), CUB-200 (Wah et al., 2011), OxfordPets-37 (Parkhi et al., 2012), StanfordCars-196 (Krause et al., 2013) and FGVC Aircraft-100 (Maji et al., 2013). For the base-novel split, we follow previous work (Khattak et al., 2023a) to divide the categories into equal halves for base and novel. For instance in CUB-200, the first 100 categories are considered base and the other 100 are considered novel and similarly for other datasets. Note that neither images nor the category names of the novel classes are seen during training.
@@
Evaluation setting and Metric. We evaluate our method under two distinct zero-shot settings. The first is the standard zero-shot base-to-novel generalization, following (Zheng et al., 2024b), where a separate model is trained on the base classes of each dataset. The second is our proposed mixed-dataset setting, designed to assess cross-domain generalization capabilities. Here, a single model is trained on a unified dataset constructed by combining the base classes from all datasets. Additionally, we also evaluate our method under few-shot classification setting with 4 shots per class. For performance measurement, we report the classification accuracy on base and novel classes, along with their Harmonic Mean (HM).
For evaluation, we use the LLM gemini-2.5-flash-lite (Comanici et al., 2025), where we provide the ground-truth category name and model predicted category name and ask the LLM if they belong to the same fine-grained category or not. This specifically helps us evaluate better for the answers where model responds a scientific name and the provided ground-truth is common name and vice-versa. We provide more details in Appendix A.1.2
Implementation Details. We use the pre-trained Qwen2.5-VL-7B-Instruct (Bai et al., 2025) as our base model and perform the RL training using the proposed method. We use K=20 during offline option generation using $K$ rollouts and other details are provided in Appendix A.1.1. For Zero Shot Base-to-Novel training, we train the model for 400 steps, whereas for mixed data training, we train it for 1 epoch. For few-shot training, we train each model for 200 steps following ViRFT. All models are trained on three A6000 GPUs with 48GB memory with an overall batch size of 6. For GRPO, a total of 4 responses are generated for each input sample. Following ViRFT, we use learning rate of $10^{-6}$, AdamW optimizer, linear scheduler and set $\lambda_{m}=\lambda_{f}=1$. For ViRFT training, we use their official code from github with some modification discussed in Appendix A.1.1.
@@ 4.2.1 Zero-Shot Base-to-Novel Generalization
We also evaluate supervised fine-tuning (SFT) as a baseline. Although SFT yields strong accuracy gains on base categories, its performance deteriorates sharply on novel categories, with an average accuracy drop of 33.5% relative to the base model and a 19.9% reduction in HM. This sharp degradation highlights SFT’s inability to generalize under the base-to-novel transfer setting. To provide a broader perspective, we also include results from proprietary models which are likely much larger. In this context, our method surpasses Grok4-fast and is on par with Gemini2.5-flash-light, though GPT-5-mini leads. We provide more results using Gemma-3 (Team et al., 2025a) base model in Appendix A.3.
@@ 4.3 Ablation Studies
Our ablation into the MCQ option generation reveals that the option construction strategy is critical for model’s performance. As detailed in Table 3, randomly selecting categories proved suboptimal, yielding only marginal gains and failing to instill robust reasoning capabilities. While employing a text embedding model from gemini embeddings (Lee et al., 2025) to generate semantically similar options offered some improvement, our proposed top-k as options proves significantly more effective. By sampling options directly from the same base model’s top-k generations achieved a substantial classification accuracy gain of 12% on base classes and 6.9% on novel classes compared to base model. This demonstrates that leveraging the base model’s own knowledge distribution to generate options is the optimal strategy for training, leading to superior generalization on both base and novel sets.
@@ A.1.2 Evaluation details
In Figure 8, we provide the prompt used for second step of multiple choice question. Here {options} refers to the options obtained from the first step of our pipeline. In Figure 9, we provide the prompt used during the evaluation. {groundtruth} refers to the groundtruth category name and {prediction} refers to the model’s predicted answer. We use "google/gemini-2.5-flash-lite-preview-06-17" as LLM for evaluation from openrouter (OpenRouter, 2024) API.